A 800 kg car travels at 25 m/s behind another car (905 kg) that travels at 15 m/s. The two cars collided and were locked together after collision. Determine the velocity of the vehicle after collision.


1
Expert's answer
2022-05-08T14:29:38-0400

Conservation of momentum says

"m_1v_1+m_2v_2=(m_1+m_2)u"

"u=\frac{m_1v_1+m_2v_2}{m_1+m_2}"

"u=\frac{800*25+905*15}{800+905}=20\:\rm m/s"